if a.,x1,x2 are in gp with common ratio r and b,y1,y2 are in gp with common ratio s and s-r=2 then the area of triangle with vertices (a''b),(x1,y1),(x2,y2) is

Grade:12

1 Answers

Vikas TU
10 years ago

a.,x1,x2 and b,y1,y2  are in gp then we can say: 2x1 = a +x2 and also 2y1 = b + y2

or, a = 2x1 - x2    and b =  2y1 - y2

Hence, area of triangle with vertices (a,b),(x1,y1),(x2,y2) is

 

Plz Approve!

Plz Approve!

Plz Approve!